Circular No. 8438

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 V5 = P/2003 YM_{159} (LINEAR-HILL) While the initial report inferred that the discovery of comet P/2003

YM_{159} at Catalina was a team discovery (thus the name ‘LINEAR-Catalina’ given on IAUC 8433), it has since been determined that observer Rik Hill was alone in discovering, measuring, and reporting the comet — thereby allowing his name to be used in place of the survey name (as also approved by the Catalina team). Consultation with the IAU Committee on Small-Body Nomenclature has yielded the decision to introduce for this comet the new principal designation P/2004 V5 and to replace the name ‘LINEAR-Catalina’ with ‘LINEAR-Hill’. The components A and B are defined as before. The following improved orbital elements from MPEC 2004-V79 are for component A, with the preliminary elements for component B being well satisfied with the same elements but with ΔT = +0.23 day.

Epoch = 2005 Mar. 11.0 TT

T = 2005 Feb. 28.6935 TT  ω = 87.^{o}6836
e = 0.445237  Ω = 47.8590 }2000.0
q = 4.410971 AU  i = 19.3582

a = 7.951091 AU n^{o} = 0.0439606 P = 22.42 years

COMET P/2004 V3 (SIDING SPRING)

Improved elliptical orbital elements from MPEC 2004-V77:

T = 2004 Nov. 13.938 TT  ω = 322.^{o}759
e = 0.45163  Ω = 356.082 }2000.0
q = 3.94084 AU  i = 50.454

a = 7.18647 AU n^{o} = 0.051160 P = 19.3 years

COMET P/2004 V4 (NEAT)

Improved elliptical orbital elements from MPEC 2004-V78:

T = 2005 Jan. 31.1328 TT  ω = 347.^{o}0816
e = 0.475536  Ω = 103.7625 }2000.0
q = 1.920247 AU  i = 12.4600

a = 3.661350 AU n^{o} = 0.1406832 P = 7.01 years

2004 November 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green